# On Supports for Graphs of Bounded Genus [^1] Parts of this work appeared in “On Hypergraph Supports” available at https://arxiv.org/abs/2303.16515, which has now been split into two papers. This is the first paper in the series, and the second paper is available at https://arxiv.org/abs/2504.05039v2

### v2

@@ 5.1 Non-blocking Chords in $abab$-free Hypergraphs
Let $C$ be a cycle embedded in the plane, and let $\mathcal{K}$ be a collection of $abab$-free subgraphs in the embedding of $C$. Then, for some disconnected $K\in\mathcal{K}$, there exists a non-blocking chord joining two disjoint runs of $K$. Further, such a chord can be computed in time $O(mn^{3})$ where $m=|\mathcal{K}|$ and $n=|C|$.
@@ 2. $(ii)$
Let $d_{0}$ denote the chord $a_{0}b_{0}$. If $d_{0}$ is non-blocking, we are done. Otherwise, if any $K_{1}\in\mathcal{K}$ is blocked by $d_{0}$, there is a vertex $k\in K_{1}$ that lies in $(b_{0},a_{0})$. Since $(b_{0},a_{0})$ does not contain a vertex of $K_{0}$, this implies $k\in K_{1}\setminus K_{0}$, and hence $d_{0}$ satisfies condition $(ii)$ of the invariant. Since we assumed the subgraphs $\mathcal{K}$ are $abab$-free, this implies that any vertex of $K_{1}$ in arc $(a_{0},b_{0})$ is contained in $K_{0}$. This ensures that condition $(i)$ of the invariant is satisfied by $d_{0}$.
@@
((a)) Ordering the vertices of $K_{0}$ in sets $A$ and $B$. Here, $A=\{a_{0},\ldots,a_{r}\}$ and $B=\{b_{0},\ldots,b_{s}\}$.
((b)) Adding chords between $A$ and $B$. If $K_{i+1}\setminus K_{0}\neq\emptyset$ in $(a_{\ell},b_{j^{\prime}})$, vertices $u,b_{j^{\prime}},v,w$ form $abab$ in $K_{i+1}$ and $K_{i}$.
@@
If $d_{i}$ is a non-blocking chord, we are done. Otherwise, let $K_{i+1}$ denote a subgraph blocked by $d_{i}$. Then, both the arcs $(a_{\ell},b_{j^{\prime}})$ and $(b_{j^{\prime}},a_{\ell})$ contain a run of $K_{i+1}$, and $a_{\ell},b_{j^{\prime}}\in K_{0}\setminus K_{i+1}$. We now show that $d_{i}$ satisfies the invariant. Most of the work will go in showing that $d_{i}$ satisfies condition $(i)$ of the invariant. We show this by contradiction - If $d_{i}$ does not satisfy invariant $(i)$, we will exhibit a pair of subgraphs violating the $abab$-free property.
@@
Finally, since $d_{i-1}$ satisfies condition $(ii)$ of the invariant, it implies that there is a vertex $w\in K_{i}\setminus K_{0}$ that lies in $(b_{j},a_{\ell})$. Note that $u\in K_{i+1}\setminus K_{0}$, $a_{\ell},b_{j^{\prime}}\in K_{0}\setminus K_{i+1}$, and $u$ lies in $(a_{\ell},b_{j^{\prime}})$. There is no vertex $x\in K_{i+1}\setminus K_{0}$ that lies in $(b_{j^{\prime}},a_{\ell})$, since otherwise we have an $abab$-pattern among $K_{i+1}$ and $K_{0}$ given by the cyclic sequence of the vertices $u,b_{j^{\prime}},x,a_{\ell}$. So, we have $K_{i+1}\subseteq K_{0}$ in $(b_{j^{\prime}},a_{\ell})$ since the arrangement is $abab$-free. However, since $(b_{j},a_{\ell})\subset(b_{j^{\prime}},a_{\ell})$ and $w\in K_{i}\setminus K_{0}$ lies in $(b_{j},a_{\ell})$, it follows that $w\not\in K_{i+1}$. Therefore, $w\in K_{i}\setminus K_{i+1}$. See Figure 4(b).
From the above arguments, it follows that the subgraphs $K_{i+1}$ and $K_{i}$ are not $abab$-free, as witnessed by the sequence of vertices $u,b_{j^{\prime}},v$ and $w$, a contradiction. Thus, $d_{i}$ satisfies condition $(i)$ of the invariant. The fact that $d_{i}$ satisfies condition $(ii)$ of the invariant follows from the fact that $K_{0}$ is minimal. Otherwise, $K_{i+1}\subseteq K_{0}$ in $(a_{\ell},b_{j^{\prime}})$ and in $(b_{j^{\prime}},a_{\ell})$, and therefore $K_{i+1}\subset K_{0}$.
@@
The chord $d=\{x,y\}$ divides the cycle $C$ into two arcs, $[x,y]$, and $[y,x]$. We construct two disjoint sub-problems on the cycles $C_{\ell}$ and $C_{r}$ obtained from $C$, where $C_{\ell}$ is obtained by adding the edge $\{x,y\}$ to the arc $[y,x]$, and $C_{r}$ is obtained by adding the edge $\{x,y\}$ to the arc $[x,y]$. The subgraphs in $C_{\ell}$ and $C_{r}$ are respectively those induced by $\mathcal{K}$, namely $\mathcal{K}_{\ell}=\{K\cap C_{\ell}:K\in\mathcal{K}\}$, and $\mathcal{K}_{r}=\{K\cap C_{r}:K\in\mathcal{K}\}$. Note that $\mathcal{K}_{\ell}$ and $\mathcal{K}_{r}$ are $abab$-free on $C_{\ell}$ and $C_{r}$, respectively. Let $n^{\ell}_{K}$ and $n^{r}_{K}$ denote respectively, the number of runs of $K\in\mathcal{K}$ in $C_{\ell}$ and in $C_{r}$. Clearly, $n^{\ell}_{K_{0}}<n_{K_{0}}$ and $n^{r}_{K_{0}}<n_{K_{0}}$. Also, for all other subgraphs $K^{\prime}\in\mathcal{K}$, $n^{\ell}_{K^{\prime}}\leq n_{K^{\prime}}$ and $n^{r}_{K^{\prime}}\leq n_{K^{\prime}}$, it follows that $\mathrm{cost}(C_{r},\mathcal{K}_{r})<\mathrm{cost}(C,\mathcal{K})$ and $\mathrm{cost}(C_{\ell},\mathcal{K}_{\ell})<\mathrm{cost}(C,\mathcal{K})$.
Hence, by the inductive hypothesis on $C_{\ell},\mathcal{K}_{\ell}$ and $C_{r},\mathcal{K}_{r}$ respectively, there exists a set of non-intersecting chords $D_{\ell}$ such that each $K\in\mathcal{K}_{\ell}$ induces a connected subgraph of $C_{\ell}\cup D_{\ell}$. Similarly, there exists a set of non-intersecting chords $D_{r}$ such that each $K\in\mathcal{K}_{r}$, induces a connected subgraph of $C_{r}\cup D_{r}$. Let $D=D_{\ell}\cup D_{r}\cup d$. If there is a $K\in\mathcal{K}$ such that $K$ induces connected subgraphs on $C_{r}\cup D_{r}$ and $C_{\ell}\cup D_{\ell}$ but $K$ is not connected on $C\cup D$, then $K\cap\{x,y\}=\emptyset$. Then, $K$ and $K_{0}$ form an $abab$-patter; a contradiction. Hence, it follows that $D$ is a set of non-intersecting chords such that each $K\in\mathcal{K}$ induces a connected subgraph of $(C\cup D)$.
@@ 6 Construction of Supports
In this section, we show that for cross-free systems on a graph of genus $g$, there exist primal, dual and intersection supports of genus at most $g$. While the existence of an intersection support implies the existence of the primal and dual supports, the construction of the intersection support uses the construction of both the primal and dual supports, and hence we start with their description first.
We obtain a polynomial time algorithm to construct a primal support when a cross-free embedding of the graph system $(G,\mathcal{H})$ is given. However, we are unable to prove a similar result for a dual support or intersection support, and this is an intriguing open question.
@@ 6.1 Primal Support
Let $(G,\mathcal{H})$ be a cross-free graph system of genus $g$ and let $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ be any $2$-coloring of the vertices of $G$. In this section, we show that we can construct a primal support $Q$ for $(G,\mathcal{H})$. If a cross-free embedding of $G$ with respect to $\mathcal{H}$ is given, then $Q$ can be computed in time polynomial in $|V(G)|$ and $|\mathcal{H}|$.
@@
Our construction of a support in the general case follows by repeated application of the Vertex Bypassing operation to a maximal red vertex. We show that each Vertex Bypassing operation decreases the number of maximal red vertices, and hence we eventually obtain a system with no maximal red vertices. While the lemma below is stated for bounded genus graphs, it works for any herediary family closed under edge-contractions.
@@
Let $(G,\mathcal{H})$ be a graph system of genus $g$ such that each $H\in\mathcal{H}$ is a connected subgraph of $G$. Then, for any coloring $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ s.t. no vertex in $\mathbf{r}(V(G))$ is maximal, there is a primal support $Q$ on $\mathbf{b}(V(G))$ of genus at most $g$. Furthermore, $Q$ can be constructed in $O(|V(G)|^{2}|\mathcal{H}|)$ time.
@@
Let $(G,\mathcal{H})$ be a graph system of genus $g$ such that each $H\in\mathcal{H}$ induces a connected subgraph of $G$, and $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ s.t. no vertex in $\mathbf{r}(V(G))$ is maximal, and $|\mathbf{r}(V(G))|=k$.
Consider an arbitrary vertex $u\in\mathbf{r}(V(G))$. Since $u$ is not maximal, by definition, there is an edge $e=\{u,v\}\in E(G)$ such that $\mathcal{H}_{e}=\mathcal{H}_{u}$. Since $\mathcal{H}_{e}\subseteq\mathcal{H}_{v}$, it follows that $\mathcal{H}_{u}\subseteq\mathcal{H}_{v}$. Let $G^{\prime}=G/\{u,v\}$, i.e., the graph obtained by contracting the edge $\{u,v\}$ with the new vertex taking the color $c(v)$. Note that $G^{\prime}$ has genus at most $g$ since contracting an edge in $G$ preserves the embedding on the same surface as $G$ does. Let $\mathcal{H}^{\prime}$ denote the subgraphs in $G^{\prime}$ corresponding to $\mathcal{H}$. Each subgraph in $\mathcal{H}^{\prime}$ clearly induces a connected subgraph of $G^{\prime}$, and $|\mathbf{r}(V(G^{\prime}))|=k-1$. By the inductive hypothesis, there is a support $Q$ for $(G^{\prime},\mathcal{H}^{\prime})$. Note that $Q$ is also a support for $(G,\mathcal{H})$ since the set of blue vertices in $H$ remains unchanged in $(G^{\prime},\mathcal{H}^{\prime})$ for each $H\in\mathcal{H}$.
Each edge contraction can be performed in time $O(|V(G)||\mathcal{H}|)$ (assuming $G$ is represented as an adjacency list) and the number of edge contractions is bounded above by $|\mathbf{r}(V(G))|\leq|V(G)|$, it follows that a primal support for $(G,\mathcal{H})$ can be computed in time $O(|V(G)|^{2}|\mathcal{H}|)$. ∎
@@
We prove by induction on the number of maximal red vertices. If no red vertex in $G$ is maximal, then by Lemma 26, there is a support $Q$ of genus $g$.
Assume the statement holds for any cross-free graph system of genus $g$ with less than $k$ maximal red vertices. Let $(G,\mathcal{H})$ be a cross-free graph system of genus $g$ and $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ a 2-coloring with $k$ maximal red vertices.
Choose an arbitrary maximal red vertex $v$ and apply $\textsc{VB}(v)$ and let $(G^{\prime},\mathcal{H}^{\prime})$ denote the resulting graph system. By Lemma 22 and Lemma 23, $(G^{\prime},\mathcal{H}^{\prime})$ is a cross-free graph system of genus at most $g$.
Let $N(v)=\{v_{0},\ldots,v_{deg(v)-1}\}$. For $i=0,\ldots,deg(v)-1$, let $u_{i}$ be the vertices added by $\textsc{VB}(v)$ subdividing edges $\{v,v_{i}\}$. Set $c(u_{i})=\mathbf{r}$, $\forall\;i=0,\ldots,deg(v)-1$. By Lemma 25, none of the $u_{i}$’s is maximal. Hence, the number of maximal red vertices in $G^{\prime}$ is at most $k-1$, and by the inductive hypothesis, there is a support $Q$ of genus at most $g$ for $(G^{\prime},\mathcal{H}^{\prime})$. It follows that $Q$ is also a primal support for $(G,\mathcal{H})$ since the set of blue vertices in each $H^{\prime}\in\mathcal{H}^{\prime}$ is identical to that in the corresponding subgraph $H\in\mathcal{H}$.
To see that $Q$ can be computed in time $O(mn^{5})$, note that the number of maximal red vertices in $G$ is at most $n$. Each vertex bypassing operation decreases the number of maximal red vertices, and hence we require at most $n$ vertex bypassing operations. By Lemma 22, each vertex bypassing operation can be done in time $O(mn^{4})$, and hence the total running time for all the vertex bypassing operations is $O(mn^{5})$. Further, note that on vertex bypassing operation at vertex $v$, $deg(v)$, many new vertices are introduced, none of which is maximal. Hence, at the end of vertex bypassing in $G$, we end up with having at most $n^{2}$ vertices such that no red vertex is maximal. By Lemma 26, it takes $O(m(n^{2})^{2})=O(mn^{4})$ time for the edge contraction operations to finally get the support $Q$. Hence, the total running time to compute $Q$ is $O(mn^{4})+O(mn^{5})=O(mn^{5})$. ∎

### v3

@@ 5.1 Non-blocking Chords in abababab-free Hypergraphs
Let $C$ be a cycle embedded in the plane, and let $\mathcal{K}$ be a collection of abababab-free subgraphs in the embedding of $C$. Then, for some disconnected $K\in\mathcal{K}$, there exists a non-blocking chord joining two disjoint runs of $K$. Further, such a chord can be computed in time $O(mn^{3})$ where $m=|\mathcal{K}|$ and $n=|C|$.
@@ 2. $(ii)$
Let $d_{0}$ denote the chord $a_{0}b_{0}$. If $d_{0}$ is non-blocking, we are done. Otherwise, if any $K_{1}\in\mathcal{K}$ is blocked by $d_{0}$, there is a vertex $k\in K_{1}$ that lies in $(b_{0},a_{0})$. Since $(b_{0},a_{0})$ does not contain a vertex of $K_{0}$, this implies $k\in K_{1}\setminus K_{0}$, and hence $d_{0}$ satisfies condition $(ii)$ of the invariant. Since we assumed the subgraphs $\mathcal{K}$ are abababab-free, this implies that any vertex of $K_{1}$ in arc $(a_{0},b_{0})$ is contained in $K_{0}$. This ensures that condition $(i)$ of the invariant is satisfied by $d_{0}$.
@@
((a)) Ordering the vertices of $K_{0}$ in sets AA and $B$. Here, $A=\{a_{0},\ldots,a_{r}\}$ and $B=\{b_{0},\ldots,b_{s}\}$.
((b)) Adding chords between $A$ and $B$. If $K_{i+1}\setminus K_{0}\neq\emptyset$ in $(a_{\ell},b_{j^{\prime}})$, vertices $u,b_{j^{\prime}},v,w$ form abababab in $K_{i+1}$ and $K_{i}$.
@@
If $d_{i}$ is a non-blocking chord, we are done. Otherwise, let $K_{i+1}$ denote a subgraph blocked by $d_{i}$. Then, both the arcs $(a_{\ell},b_{j^{\prime}})$ and $(b_{j^{\prime}},a_{\ell})$ contain a run of $K_{i+1}$, and $a_{\ell},b_{j^{\prime}}\in K_{0}\setminus K_{i+1}$. We now show that $d_{i}$ satisfies the invariant. Most of the work will go in showing that $d_{i}$ satisfies condition $(i)$ of the invariant. We show this by contradiction - If $d_{i}$ does not satisfy invariant $(i)$, we will exhibit a pair of subgraphs violating the abababab-free property.
@@
Finally, since $d_{i-1}$ satisfies condition $(ii)$ of the invariant, it implies that there is a vertex $w\in K_{i}\setminus K_{0}$ that lies in $(b_{j},a_{\ell})$. Note that $u\in K_{i+1}\setminus K_{0}$, $a_{\ell},b_{j^{\prime}}\in K_{0}\setminus K_{i+1}$, and $u$ lies in $(a_{\ell},b_{j^{\prime}})$. There is no vertex $x\in K_{i+1}\setminus K_{0}$ that lies in $(b_{j^{\prime}},a_{\ell})$, since otherwise we have an abababab-pattern among $K_{i+1}$ and $K_{0}$ given by the cyclic sequence of the vertices $u,b_{j^{\prime}},x,a_{\ell}$. So, we have $K_{i+1}\subseteq K_{0}$ in $(b_{j^{\prime}},a_{\ell})$ since the arrangement is abababab-free. However, since $(b_{j},a_{\ell})\subset(b_{j^{\prime}},a_{\ell})$ and $w\in K_{i}\setminus K_{0}$ lies in $(b_{j},a_{\ell})$, it follows that $w\not\in K_{i+1}$. Therefore, $w\in K_{i}\setminus K_{i+1}$. See Figure 4(b).
From the above arguments, it follows that the subgraphs $K_{i+1}$ and $K_{i}$ are not abababab-free, as witnessed by the sequence of vertices $u,b_{j^{\prime}},v$ and $w$, a contradiction. Thus, $d_{i}$ satisfies condition $(i)$ of the invariant. The fact that $d_{i}$ satisfies condition $(ii)$ of the invariant follows from the fact that $K_{0}$ is minimal. Otherwise, $K_{i+1}\subseteq K_{0}$ in $(a_{\ell},b_{j^{\prime}})$ and in $(b_{j^{\prime}},a_{\ell})$, and therefore $K_{i+1}\subset K_{0}$.
@@
The chord $d=\{x,y\}$ divides the cycle $C$ into two arcs, $[x,y]$, and $[y,x]$. We construct two disjoint sub-problems on the cycles $C_{\ell}$ and $C_{r}$ obtained from $C$, where $C_{\ell}$ is obtained by adding the edge $\{x,y\}$ to the arc $[y,x]$, and $C_{r}$ is obtained by adding the edge $\{x,y\}$ to the arc $[x,y]$. The subgraphs in $C_{\ell}$ and $C_{r}$ are respectively those induced by $\mathcal{K}$, namely $\mathcal{K}_{\ell}=\{K\cap C_{\ell}:K\in\mathcal{K}\}$, and $\mathcal{K}_{r}=\{K\cap C_{r}:K\in\mathcal{K}\}$. Note that $\mathcal{K}_{\ell}$ and $\mathcal{K}_{r}$ are abababab-free on $C_{\ell}$ and $C_{r}$, respectively. Let $n^{\ell}_{K}$ and $n^{r}_{K}$ denote respectively, the number of runs of $K\in\mathcal{K}$ in $C_{\ell}$ and in $C_{r}$. Clearly, $n^{\ell}_{K_{0}}<n_{K_{0}}$ and $n^{r}_{K_{0}}<n_{K_{0}}$. Also, for all other subgraphs $K^{\prime}\in\mathcal{K}$, $n^{\ell}_{K^{\prime}}\leq n_{K^{\prime}}$ and $n^{r}_{K^{\prime}}\leq n_{K^{\prime}}$, it follows that $\mathrm{cost}(C_{r},\mathcal{K}_{r})<\mathrm{cost}(C,\mathcal{K})$ and $\mathrm{cost}(C_{\ell},\mathcal{K}_{\ell})<\mathrm{cost}(C,\mathcal{K})$.
Hence, by the inductive hypothesis on the pair $(C_{\ell},\mathcal{K}_{\ell})$, there exists a set of non-intersecting chords $D_{\ell}$ such that each $K\in\mathcal{K}_{\ell}$ induces a connected subgraph of $C_{\ell}\cup D_{\ell}$. Similarly, there exists a set of non-intersecting chords $D_{r}$ such that each $K\in\mathcal{K}_{r}$, induces a connected subgraph of $C_{r}\cup D_{r}$. Let $D=D_{\ell}\cup D_{r}\cup d$. If there is a subgraph $K\in\mathcal{K}$ such that $K$ induces connected subgraphs on both $C_{r}\cup D_{r}$ and $C_{\ell}\cup D_{\ell}$ but $K$ is not connected on $C\cup D$, then $K\cap\{x,y\}=\emptyset$. It follows that $K$ is blocked by $d$; a contradiction that $d$ is a non-blocking chord. Hence, it $D$ is a set of non-intersecting chords such that each $K\in\mathcal{K}$ induces a connected subgraph of $(C\cup D)$.
@@ 6 Construction of Supports
In this section, we show that for cross-free systems on a graph of genus $g$, there exist primal, dual and intersection supports of genus at most $g$. While the existence of an intersection support implies the existence of the primal and dual supports, the construction of the intersection support uses the construction of both the primal and dual supports, and hence, we start with their description first.
@@ 6.1 Primal Support
Let $(G,\mathcal{H})$ be a cross-free graph system of genus $g$ and let $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ be any $2$-coloring of the vertices of $G$. In this section, we show that we can construct a primal support $Q$ for $(G,\mathcal{H})$ such that the genus of $Q$ is at most $g$.
@@
Our construction of a support in the general case follows by repeated application of the Vertex Bypassing operation to a maximal red vertex. We show that each Vertex Bypassing operation decreases the number of maximal red vertices, and hence we eventually obtain a system with no maximal red vertices. While the lemma below is stated for bounded genus graphs, it works for any hereditary family closed under edge-contractions.
Before stating the lemma, we introduce the following terminology. A pair of vertices $u,v\in V(G)$ are said to be twins if $\mathcal{H}_{u}=\mathcal{H}_{v}$. We say that $u$ and $v$ are adjacent twins if in addition $u$ and $v$ are adjacent vertices in $G$. If $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ is a 2-coloring of $V(G)$ and $u,v\in\mathbf{r}(V(G))$ are twins, we say that $u$ and $v$ are red twins.
@@
Let $(G,\mathcal{H})$ be a graph system of genus $g$ such that each $H\in\mathcal{H}$ is a connected subgraph of $G$. Let $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ be any 2-coloring of $V(G)$ such that no vertex in $\mathbf{r}(V(G))$ is maximal and there are no adjacent twins $u,v\in\mathbf{r}(V(G))$.
@@
Let $(G,\mathcal{H})$ be a graph system with a 2-coloring $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ satisfying the conditions of the lemma such that $|\mathbf{r}(V(G))|=k$.
For each vertex $u\in\mathbf{r}(V(G))$, we do the following: Since $u$ is not maximal, by definition, there is an edge $e=\{u,v\}\in E(G)$ such that $\mathcal{H}_{e}=\mathcal{H}_{u}$. We chose one such edge $e$ incident to $u$, breaking ties arbitrarily. Since $\mathcal{H}_{e}\subseteq\mathcal{H}_{v}$, it follows that $\mathcal{H}_{u}\subseteq\mathcal{H}_{v}$. We orient this edge from $u$ to $v$. Further, if $v\in\mathbf{r}(V(G))$, then $\mathcal{H}_{u}\subset\mathcal{H}_{v}$ since there are no adjacent twins in $\mathbf{r}(V(G))$.
It follows that the set of oriented edges induces a directed acyclic subgraph. Further, since there are no twins in $\mathbf{r}(V(G))$, this directed acyclic graph is an oriented forest. Since no red vertex is maximal, it implies that each tree in the forest has its edges oriented towards the root, and the root must be a blue vertex. Choose an arbitrary tree in the forest and choose an edge $\{u,v\}$ within the tree, where $v$ is the root of the tree. Let $G^{\prime}=G/\{u,v\}$, i.e., the graph obtained by contracting the edge $\{u,v\}$ with the new vertex taking the color $c(v)=\mathbf{b}$, since $v$, the root of the tree is colored $\mathbf{b}$. Let $\mathcal{H}^{\prime}$ be the resulting family of subgraphs, where for each $H\in\mathcal{H}$, we have a subgraph $H^{\prime}\in\mathcal{H}^{\prime}$ of $G^{\prime}$ induced on the vertices $V(H)\cap V(G^{\prime})$.
Note that no red vertex is maximal in $G^{\prime}$ since each red vertex in $G^{\prime}$ has an outgoing edge incident on it. Further, there are no adjacent twins in $G^{\prime}$ among the red vertices as the edge contraction did not introduce new adjacencies between the red vertices. To see this note that the vertex obtained by contracting the edge $\{u,v\}$ is colored $\mathbf{b}$, and hence, the adjacency among the red vertices is unchanged. Since there are no adjacent red twins in $G$ it implies that there are no adjacent twins in $G^{\prime}$. Finally, note that $G^{\prime}$ has genus at most $g$ since contracting an edge in $G$ preserves the embedding on the same surface as $G$ does.
Since $G^{\prime}$ has k1k-1 red vertices, by the inductive hypothesis, there is a primal support $Q$ on the blue vertices of $G^{\prime}$. The graph $Q$ is also a support for $(G,\mathcal{H})$ since the set of blue vertices in $H$ remains unchanged in $(G^{\prime},\mathcal{H}^{\prime})$ for each $H\in\mathcal{H}$. ∎
Now we prove that given a cross-free graph system $(G,\mathcal{H})$ of genus $g$ and a 2-coloring $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$, there is a primal support of genus at most $g$. To establish the result, we repeatedly apply edge contractions for adjacent red twins, and use an induction argument on the maximum depth of a maximal red vertex via the applications of the vertex bypassing operation in combination with Lemma 25.
@@
If there are adjacent red twins $u$ and $v$, we contract the edge $\{u,v\}$ to a single vertex and assign it color $\mathbf{r}$. This preserves the cross-freeness among subgraphs in $\mathcal{H}$. Since for any $H\in\mathcal{H}$, the set of vertices in $\mathbf{b}(V(G))\cap V(H)$ remains unchanged, a primal support for the modified system is also a support for the given system $(G,\mathcal{H})$. So, we assume wlog that there is no adjacent red twin.
We prove the result by induction on the maximum depth of a maximal red vertex. Let $d$ be the maximum depth of a maximal red vertex. Note that a non-maximal red vertex can have depth more than $d$. For $i\in\{1,2,\ldots,d\}$, let $S_{i}\subseteq\mathbf{r}(V(G))$ be the set of maximal red vertices of depth $i$. Let $d=1$. It follows from the definition of a maximal vertex that for each $v\in S_{d}$ the subgraph containing $v$ consists of a single vertex $v$. We remove all such subgraphs from $\mathcal{H}$, since they do not contain any blue vertex and thus the conditions of support are trivially satisfied for them. Let $\mathcal{H}_{1}=\mathcal{H}\setminus S_{1}$. Then, the graph system $(G,\mathcal{H}_{1})$ satisfies the conditions of Lemma 26, and thus admits a primal support.
Assume the statement holds for any cross-free graph system of genus $g$ with $d<k$, for some positive integer $k\geq 2$. Let $(G,\mathcal{H})$ be a cross-free graph system of genus $g$ and $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ a 2-coloring such that $d=k$, i.e., the maximum depth of a maximal red vertex is $k$. We apply the process below to obtain a cross-free system $(G^{\prime},\mathcal{H}^{\prime})$ such that the maximum depth of a maximal red vertex is at most k1k-1, and there are no adjacent twins among the red vertices. We do this in two steps.
In the first step, we apply vertex bypassing to all the vertices in $S_{k}$, and we assign color $\mathbf{r}$ to all the newly added vertices during the vertex bypassing process. Note that if $v\in S_{k}$, then by Lemma 25 none of the newly added vertices is maximal. By Lemma 23, the resulting graph system remains cross-free. In the second step, we repeatedly contract all the adjacent red twins and color the contracted vertex $\mathbf{r}$. Clearly, the resulting graph system $(G^{\prime},\mathcal{H}^{\prime})$ remains cross-free and does not contain any adjacent red twin.
In $(G^{\prime},\mathcal{H}^{\prime})$, we have $d\leq k-1$. To see this, note that all the vertices of $G$ in $S_{k}$ have been bypassed. It is possible that during the edge contractions in the second step above, two non-maximal adjacent red twins $x$ and $y$ may result in a maximal red vertex $z$. However, at least one of $x$ and $y$ must be added during the vertex bypassing operation since we assumed that $G$ does not contain adjacent red twins. By Lemma 25, $0pt(z)\leq k-1$.
Therefore, $(G^{\prime},\mathcal{H}^{\prime})$ is a cross-free graph system of genus $g$ such that the maximum depth of a maximal red vertex is at most k1k-1, and there are no adjacent red twins. By the induction hypothesis, $(G^{\prime},\mathcal{H}^{\prime})$ admits a primal support $Q$ of genus at most $g$, which is also a primal support for $(G,\mathcal{H})$ since the set of blue vertices remains unchanged for each subgraph in $\mathcal{H}$.